# Online Navigation Planning for Long-term Autonomous Operation of Underwater Gliders

### v1

@@ VII-B Planner Evaluation in Simulation
Figure 9: Transects completed by the proposed GLAE system in the MOGli field deployment using the Planner and the Straight-to-Goal baseline. The Planner results in transects with shorter minimum, mean, and maximum durations. Transects produced by the two methods are also discernible visually. The transects completed by the Planner have statistically significantly shorter distances and avoid overshooting the western goal.
Table IV shows the results in simulation comparing the Straight-to-Goal and Planner algorithms. The Planner improves over Straight-to-Goal in duration and transect length for all deployments and scenarios, with gains of up to $9.88\%$ and $16.51\%$ respectively. For MOGli, the gains in duration over the baseline are small in an absolute sense at slightly over $30$ minutes per transect, equating to a $2.39\%$ decrease. This region does not exhibit strong currents, as corroborated by the small differences in the performance of STG between the most Favourable and Unfavourable scenarios. Therefore, our currents-aware approach does not have much room for improvement over this simple algorithm. The gains are more pronounced for eSWEETS3, where the Planner leads to decreases of $9.88\%$ and $8.92\%$ respectively in dive duration. The reductions in transect length (of $4.23\%,16.51\%,13.94\%$ for the three deployments) are more pronounced than the duration reductions. This is an important secondary metric as sampling along trajectories that are as straight as possible is key to many scientific objectives.
VII-C Planner Evaluation in Field Deployments
TABLE V: Results comparing the Straight-to-Goal and Planner algorithms in the field deployments. Bold denotes the better-performing algorithm for each deployment and metric.
Deployment Algorithm Mean Duration (hrs) Min Duration (hrs) Max Duration (hrs) Mean Number Dives Min Number Dives Max Number Dives Mean Transect Length (km)
@@ VII-C Planner Evaluation in Field Deployments
Figure 10: Proportions of actions chosen by the Planner with respect to the bearing to goal during each of the three field deployments. The stronger currents of the eSWEETS3 region resulted in a wider distribution of frequencies of chosen actions.
@@
Results for the real-world deployments are reported in Table V, showing the Planner outperformed the baseline in $2$ out of $3$ deployments. For MOGli, as in simulation, the Planner leads to a reduction in the mean duration of approximately $30$ minutes on average; the minimum and maximum durations across all transects are also reduced. The Planner yields a reduction in the transect length that is statistically significant at the 5% level ($p=0.041$ and rank-biserial $r=0.376$ under the nonparametric Mann-Whitney U test with $20$ samples for STG and $21$ for the Planner). However, given the number of transects recorded (see Table II), statistical significance was not obtained for other deployments and metrics. These field results should be treated as corroborating evidence alongside the larger-scale simulation study rather than as standalone proof of generality.
The Planner also performed better than the baseline for eSWEETS3 North in all metrics. For eSWEETS3 South, the mean values for the metrics were slightly worse, while the minimum / maximum ranges still matched or improved over STG. The advantage of the Planner in eSWEETS3 North is more substantial than that of STG in eSWEETS3 South.
@@
In Figure 10, we analyse the frequency of the actions taken by the Planner in the three deployments. While the STG action ($\alpha=0$) is the most selected overall, the Planner chooses to steer frequently. In the eSWEETS3 region, which is characterised by stronger tidal currents, the distribution of chosen actions is wider and flatter, and even the most extreme actions are still chosen $15\%$ of the time in eSWEETS3 North. Given this, the Planner could have benefitted from an even wider set of allowed actions in eSWEETS3 North.
Regarding the alignment between simulation and real-world results, for MOGli, the duration and number of dives are mostly aligned, while the transect lengths are noticeably shorter in simulation. This is consistent with the distance underestimation observed in the simulator evaluation.
@@
For the eSWEETS3 deployments, values for all metrics were consistently higher in simulation compared to the field deployments. Given the values reported in Section VII-A, simulator predictions of dive distance and durations are well-aligned with field observations. Consequently, the difference in transect metrics is attributable to the goal achievement conditions in the deployed autonomous system being broader (3 conditions, as described in Sections V-C and VI-A) than those in simulation (1 condition). Requiring surfacing within a tight radius in simulation will result in overshooting the goal and hence longer transects until the goal is eventually achieved. Concretely, the goal was deemed achieved for eSWEETS3 by reaching it underwater in $43\%$ of cases, having traversed 90% of the transect prior to safety box re-entry in $22\%$ of cases, and by surfacing within the goal radius in only $35\%$ of transects.
VIII Discussion
@@
We highlight the following challenges and areas for future work. We aim to generalise our method to also consider the optimisation of energy use, which is feasible with the adopted formulations. As the duration and energy use objectives conflict, a multi-objective approach [42] is required. Extending the method to incorporate constraints and guarantees, such as staying within a specified survey area or avoiding certain areas (e.g., shipping lanes) would broaden the applicability of the method. The problem may also be framed as a POMDP [43] in order to maintain beliefs about the nature of forecast errors, which requires accurate in-situ current measurements.
@@
Replacing or augmenting the simulator with a learned model, which has proven successful with a variety of planning and reinforcement learning approaches [44, 45], may speed up execution and thus enable planning over a longer horizon. As shown by our simulator evaluation, further improvements in accuracy are possible. Fitting simulator behaviours using different vehicles and deployment environments is a significant challenge due to distribution shifts. Further developments to the planner can include the aggregation of continuous actions for root-parallel MCTS [46], improving the exploration strategy [47], and contextually deciding which actions should be prioritised.
@@
Evaluating glider navigation planning algorithms in the field is inherently challenging given the long transect traversals and, consequently, the low number of samples obtained. If resources permit, running algorithms on gliders deployed side-by-side in a paired experimental design would reduce variability and temporal dependencies compared to the alternative-control design we adopted. However, replaying on-field conditions in suitably calibrated simulators, as performed in our evaluation, is ultimately necessary for statistically robust findings.
@@
Failures in transferring instructions also pose a unique challenge, as they can cause the glider to turn around towards a stale waypoint. Such failures disproportionately impact navigation plans with waypoints that steer away from the goal to compensate for currents. This phenomenon highlights the trade-off between the gains of online planning for steering given up-to-date current forecasts versus the negative impacts of using a stale online plan due to failed instruction transfers. This can be addressed by explicitly reasoning about the time available for planning [48] and limiting it in order to improve the probability of successful transfer.
@@
In terms of applications, we envisage leveraging the present work as a building block. Analysing ocean fronts for features such as temperature [49] or algal blooms [50] can be achieved by creating a higher-level component that reasons about where the highest-value observation(s) can be gathered and tasking the glider accordingly. For going beyond the single-glider scenario, coordination strategies can be considered in a multi-robot setting.
IX Conclusion
@@
In this paper, we have considered the problem of navigation planning for underwater gliders, an invaluable tool for ocean sampling. We have taken important steps for addressing methodological and systems gaps towards achieving long-term autonomous glider operation. Our work proposed a glider simulator and a procedure for adjusting its parameters using historical dives performed in the real world. We have also contributed a sample-based online planning technique for deciding how to control the robot with the objective of minimising duration to reach the goal. Lastly, in collaboration with pilots and domain experts, we have designed the Glider Long-term Autonomy Engine system, which is suitable for long-term missions. We have thoroughly validated our approach in simulation and through extensive trials in the North Sea. Our work paves the way towards autonomous management of fleets of gliders and can ultimately support advancements in marine science and policy.
@@ Acknowledgments
This work was supported by the UK Natural Environment Research Council (NERC) under the UKRI TWINE programme (grant number NE/Z503381/1) and the Doctoral Training Partnership in Environmental Research (NE/S007474/1), the UK Engineering and Physical Sciences Research Council (EPSRC) From Sensing to Collaboration Programme Grant (EP/V000748/1), and the Innovate UK AutoInspect Grant (1004416). The authors would like to gratefully acknowledge the contribution of the Met Office and the University of East Anglia, who gave permission for the GLAE system to be trialled in their glider campaigns. We thank Christopher Auckland & Rob Hall (University of East Anglia), who helped design the additional features for the eSWEETS3 campaign. Lastly, we acknowledge the use of the University of Oxford Advanced Research Computing (ARC) facility in carrying out this work (see http://dx.doi.org/10.5281/zenodo.22558).

### v2

@@ 7.B Planner Evaluation in Simulation
Figure 12: Transects completed by the proposed GALE system in the MOGli field deployment using the Planner and the Straight-to-Goal baseline. Only transects in which all transfers of control instructions succeeded are shown. The goal radii are represented by two dotted circles around the goals. The Planner results in transects with shorter minimum, mean, and maximum durations. Transects produced by the two methods are also discernible visually. The transects completed by the Planner have statistically significantly shorter path lengths and avoid overshooting the western goal.
Table 4 shows the results comparing the Straight-to-Goal and Planner algorithms in simulation. The Planner improves over Straight-to-Goal in duration and path length in all settings tested. The most significant improvement is obtained for the Favourable scenario in the eSWEETS3 North deployment, in which the planner reduces the mean transect duration by $7$ hours and the path length by $11$ km. For MOGli, the reductions in duration over the baseline are small in an absolute sense at slightly over $30$ minutes per transect, equating to a $2.39\%$ decrease. This region does not exhibit strong currents, as corroborated by the small differences in the performance of STG between the most Favourable and Unfavourable scenarios. Therefore, our currents-aware approach does not have much room for improvement over the STG algorithm. The improvements are more pronounced for eSWEETS3, where the Planner leads to decreases of $9.88\%$ and $8.92\%$ respectively in dive duration. The reductions in path length of $4.23-13.94\%$ are more pronounced than the duration reductions. This is an important secondary metric as sampling along trajectories that are as straight as possible facilitates robust environmental data analysis and helps achieve many scientific objectives.
7.C Planner Evaluation in Field Deployments
Table 5: Results comparing the Straight-to-Goal and Planner algorithms in the field deployments. Bold denotes the better-performing algorithm for each deployment and metric.
Deployment Algorithm Mean Duration (hrs) Min Duration (hrs) Max Duration (hrs) Mean Number Dives Min Number Dives Max Number Dives Mean Path Length (km)
@@ 7.C Planner Evaluation in Field Deployments
Results for the real-world deployments are reported in Table 5. These show that the Planner outperformed the baseline in $2$ out of $3$ deployments. For MOGli, as in simulation, the Planner leads to a reduction in the mean duration of approximately $30$ minutes on average; the minimum and maximum durations across all transects are also reduced. The Planner yields a reduction in the path length that is statistically significant at the 5% level ($p=0.041$ and rank-biserial $r=0.376$ under the nonparametric Mann-Whitney U test with $20$ samples for STG and $21$ for the Planner). However, given the number of transects recorded (see Table 2), statistical significance was not obtained for other deployments and metrics. These field results should be treated as corroborating evidence alongside the larger-scale simulation study rather than as standalone proof of generality. The Planner also performed better than STG for eSWEETS3 North in all metrics. For eSWEETS3 South, the mean values were slightly worse, while the minimum / maximum ranges still matched or improved over STG.
@@
In Figure 13, we analyse the frequency of the actions taken by the Planner in the three deployments. While the STG action ($\alpha=0$) is the most selected overall, the Planner often chooses to steer relative to the goal. In the eSWEETS3 region, characterised by stronger tidal currents, the distribution of chosen actions is wider and flatter, and even the most extreme actions are still chosen $15\%$ of the time in eSWEETS3 North. The strong tidal currents in the eSWEETS3 region also resulted in surfacings outside of the safety polygon. 27 such surfacings (20.3%) occurred for eSWEETS3 North and 4 (2.9%) for eSWEETS3 South.
Figure 13: Proportions of actions chosen by the Planner during the three field deployments, split between the East and West directions of travel. The stronger currents of the eSWEETS3 region resulted in a wider distribution of action frequencies.
7.D Alignment of Simulation and Real-World Results
@@
For MOGli, the duration and number of dives are mostly aligned, while the path lengths are noticeably shorter in simulation. This is consistent with the distance underestimation observed in the simulator evaluation. For the eSWEETS3 deployments, values for all metrics were consistently higher in simulation compared to the field deployments. Given the values reported in Section 7.A, simulator predictions of dive distance and durations are well-aligned with field observations. Consequently, the difference in transect metrics is attributable to the goal achievement conditions in the deployed autonomous system being broader (3 conditions, as described in Sections 5.A and 6.A) than those in simulation (1 condition). Requiring surfacing within a tight radius in simulation will result in overshooting the goal and hence longer transects until the goal is eventually achieved. Concretely, in the eSWEETS3 field trials, the goal was deemed achieved by reaching it underwater in $46\%$ of cases, having traversed 90% of the transect length prior to safety polygon re-entry in $22\%$ of cases, and by surfacing within the goal radius $\rho$ in only $32\%$ of transects.
7.E Summary of Results
Overall, our results demonstrate that the proposed online planning techniques outperform the standard Straight-to-Goal baseline. Improvements of up to 9.88% in dive duration and 16.51% in path lengths were obtained in simulations calibrated on real-world data. Our system also obtained a statistically significant path length reduction of 9.55% in a field deployment. Compared to prior studies that are limited to simulation or short-term field trials, these results provide evidence of performance gains in real-world operational campaigns. A key achievement of our work is the demonstration of sustained autonomous operation at scale. The GALE system was deployed in two operational campaigns totalling over 3 months and 1000 km of autonomous glider operation, successfully addressing the challenges of environmental and actuation uncertainty, intermittent communication, and long-horizon navigation planning.
8 DISCUSSION
8.A Lessons Learned and Engineering Considerations
The feedback of pilots involved in the two campaigns highlighted the value of the system in automating standard operation. The system reduced manual effort and allowed pilots to focus on monitoring and non-routine interventions for other campaigns. In the eSWEETS3 campaign, the safety polygon feature allowed for automated intervention when pilot supervision may not have been available otherwise. Additionally, the straighter transects attained by our planning approach lead to more consistent sampling along predefined paths, which supports robust environmental data analysis and aligns with common scientific objectives.
The early stages of the project built gradually towards achieving the full control loop. For testing the systems integration, a “shadow” glider was initially configured. This glider duplicated the status updates received from the real glider. The control instructions generated by GALE were placed in a special folder where they could be sanity-checked by pilots but not acted upon. Subsequently, pilots began transferring the validated control instructions manually to the real glider. Once confidence in the system was achieved, fully autonomous piloting by GALE was enabled during working hours, with manual piloting outside of this window. By the start of the MOGli deployment, pilots were comfortable with simply monitoring the operation of the fully autonomous system regularly.
Regarding the generality of our system, given the modular architecture, GALE is extensible to other glider designs such as the Seaglider [45]. From a software engineering standpoint, this would require replacing the SFMC and SFMC Adapter with vehicle-specific piloting systems and adapters. From a methodological point of view, while Seagliders have slightly different low-level actuators, they also require the specification of waypoint and control parameters in an online piloting loop. Thus, our methodology is applicable to other glider designs with minimal modifications.
8.B Challenges and Future Work
@@
We aim to generalise our method to also consider the optimisation of energy use, which is a feasible extension of the proposed formulation. As the duration and energy use objectives conflict, a multi-objective approach [46] is required. This would allow the selection of more energy-intensive controls during periods with strong currents in order to make sufficient progress towards the scientific objectives. Extending the method to incorporate constraints and guarantees, such as staying within a specified survey area or avoiding certain areas (e.g., shipping lanes) would broaden the applicability of the method. The problem may also be framed as a POMDP [47] in order to maintain beliefs about the nature of forecast errors, which would require in-situ current measurements.
@@
Replacing or augmenting the simulator with a learned model, which has proven successful with a variety of planning and reinforcement learning approaches [48, 49], may speed up execution and thus enable planning over a longer horizon. As shown by our simulator evaluation, further improvements in accuracy are possible. However, fitting simulator behaviours using different vehicles and deployment environments is a significant challenge due to distribution shifts. Further developments to the planner can include the aggregation of continuous actions for root-parallel MCTS [50], improving the exploration strategy [51], and contextually deciding which actions to prioritise.
@@
Evaluating glider navigation planning algorithms in the field is inherently challenging given the long traversals and, consequently, the low number of transects. If resources permit, running algorithms on gliders deployed side-by-side in a paired experimental design would reduce variability and temporal dependencies compared to the alternative-control design we adopted. Regardless, replaying on-field conditions in suitably calibrated simulators, as performed in our evaluation, is ultimately necessary for statistically robust findings.
@@
Failures in transferring instructions to the glider also pose a unique challenge, as they can cause the glider to turn around towards a stale waypoint. Such failures disproportionately impact navigation plans with waypoints that steer away from the goal to compensate for currents. This phenomenon highlights the trade-off between the gains of online planning for steering given up-to-date current forecasts versus the negative impacts of using a stale online plan due to failed instruction transfers. This can be addressed by explicitly reasoning about the time available for planning [52] and limiting it in order to improve the probability of instruction transfers succeeding.
@@
In terms of applications, we envisage leveraging the present work as a building block. The system can operate in regions with high eddy activity using high-resolution ocean forecast models, a task that is challenging for current piloting methods. Identifying ocean fronts from horizontal temperature gradients [53] or tracking algal blooms [54] can be achieved by creating a higher-level component that reasons about where the highest-value observation(s) can be gathered and tasking the glider accordingly. Decision-making could also be implemented to minimise biases identified in existing sampling strategies (e.g., [55]). For going beyond the single-glider scenario, coordination strategies can be considered in a multi-robot setting. High-level methods can be constructed that assign goals or regions to individual gliders, with our planner being used to compute uncertainty-aware navigation plans for each vehicle.
9 CONCLUSION
@@
In this paper, we have considered the problem of navigation planning for underwater gliders, a type of robot that has become invaluable for ocean sampling. We have taken important steps to address the current methodological and systems gaps towards achieving long-term autonomous glider operation. Our work proposed a glider simulator and a procedure for adjusting its parameters using historical dives performed in the real world. We have also contributed a sample-based online planning technique for deciding how to control the glider with the objective of minimising duration to reach the goal. Lastly, in collaboration with pilots and domain experts, we have designed the Glider Autonomy Long-term Planning Engine, which is suitable for long-term and multi-platform missions.
We have thoroughly validated our approach in simulation and through operational campaigns in the North Sea, in which our system was used to control gliders collecting data towards refining weather models and understanding the impacts of offshore wind farms. Our work improves the sampling of scientific data and paves the way towards the autonomous management of fleets of gliders at scale, reducing costs and allowing human pilots to focus on monitoring and non-routine interventions. Our contributions ultimately support advancements in marine science and policy.
@@ ACKNOWLEDGMENTS
This work was supported by the UK Natural Environment Research Council (NERC) under the UKRI TWINE programme (grant number NE/Z503381/1) and the Doctoral Training Partnership in Environmental Research (NE/S007474/1). It was also supported by the UK Engineering and Physical Sciences Research Council (EPSRC) From Sensing to Collaboration Programme Grant (EP/V000748/1) and the Innovate UK AutoInspect Grant (1004416). The authors would like to gratefully acknowledge the contribution of the UK Met Office and the University of East Anglia, who gave permission for the GALE system to be trialled in their glider campaigns. The eSWEETS3 glider campaign was funded by the UK NERC (grant number NE/X004872/1). Lastly, we acknowledge the use of the University of Oxford Advanced Research Computing (ARC) facility in carrying out this work (see http://dx.doi.org/10.5281/zenodo.22558).